There are 3 black and 4 white balls in a box. One ball is taken out at random. If it is black, then two white balls are put back in the box. If it is white, then one black ball is put back in the box. After that procedure another ball is taken out of the box. What is the probability that the first ball taken was white is the second ball taken was white?
Answer #1

There are 3 black and 4 white balls in a box

P[ the ball in first draw is black ] = 3/7

P[ the ball in first draw is white ] = 4/7

Case 1 : If black ball was drawn in the first draw

If it is black, then two white balls are put back in the box

New composition : There are 3 black and 6 white balls in a box

P[ drawing white ball in second draw | first ball was black ] = 6/9 = 2/3

Case 2 : If white ball was drawn in the first draw

If it is white, then one black ball is put back in the box

New composition : There are 4 black and 4 white balls in a box

P[ drawing white ball in second draw | first ball was white ] = 4/8 = 1/2

P[ drawing white ball in second draw ] = P[ drawing white ball in second draw | first ball was white ]*P[ the ball in first draw is white ] + P[ drawing white ball in second draw | first ball was black ]*P[ the ball in first draw is black ]

P[ drawing white ball in second draw ] = (4/7)(1/2) + (3/7)(2/3)

P[ drawing white ball in second draw ] = 2/7 + 2/7

P[ drawing white ball in second draw ] = 4/7

P[ both balls are white ] = P[ the ball in first draw is white ]*P[ drawing white ball in second draw ]

P[ both balls are white ] = (4/7)(4/7)

P[ both balls are white ] = 16/49